The accountant for Perezs Dress Shop prepared the following cash budget. Perezs desires to maintain a cash cushion of $15,0Req A Req B and C Complete the cash budget by filling in the missing amounts. (Any shortages or repayments should be indicateReq A Req B and C Determine the amount of net cash flows from both operating and financing activities Perezs will report on
Answer #1

Solution a:  

B Perezs Dress shop Cash Budget July August September $ 48,000 $191,000 $239,000 $15,000 $211,000 $226,000 $15,000 $251,600

Solution b:

Net cash (operating activities) = Total cash receipts - Total cash disbursements

= (191000+211000+251600) - (231026+211860+246663)

= - $35,949 (Outflow)

Solution c:

Net cash (financing activities) = 7026+860 - 4937 = $2949 (Inflow)